# Large Angular Momentum

### v1

@@ Abstract
Quantum states of a spin $\tfrac{1}{2}$ (a qubit) are parametrized by the space ${\mathbf{CP}}^{1}\sim S^{2}$, the Bloch sphere. A spin $j$ (a $12j+$ 1-state system) for generic $j$ is represented instead by a point of a larger space, ${\mathbf{CP}}^{2j}$. Here we study the angular momentum/spin in the limit, $j\to\infty$. The state, $({\hat{\mathbf{J}}}\cdot{\mathbf{n}})\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle$, where ${\hat{\mathbf{J}}}$ is the angular momentum operator and ${\mathbf{n}}$ stands for a generic unit vector in ${\bf R}^{3}$, is found to behave as a classical angular momentum, $j\,{\mathbf{n}}$. We discuss this phenomenon, by analysing the Stern-Gerlach experiments, the angular-momentum composition rule, and the rotation matrix. This problem arose from the consideration of a macroscopic body under an inhomogeneous magnetic field. Our observations help to explain how classical mechanics (with unique particle trajectories) emerges naturally from quantum mechanics in this context, and at the same time, make the widespread notion that large spins somehow become classical, a more precise one.
@@ 1 Introduction
It is a widely shared view that in the limit of large spin (angular momentum) the quantum-mechanical spin somehow becomes classical. After all, the large spin limit ($S/\hbar\to\infty$) is equivalent to the $\hbar\to 0$ limit, and, in accordance with the general idea of the semi-classical limit, physics of large angular momentum is expected to become that of classical angular momentum. But to the best of the authors’ knowledge, it has not been clearly exhibited in the literature how exactly this works  [^1] Understandably, there is a vast literature about large angular momentum, from the early days of quantum mechanics to today, in a wide range of research fields, in various contexts and problems, and also based on different motivations. In order to avoid being biased and necessarily incomplete, we shall refrain in this work from attempting to cite or comment on any of them.. It is the aim of this note to fill in this gap.
The question came under a sharp focus in a general discussion on emergence of classical mechanics for a macroscopic body from quantum mechanics [1, 2]. Knowing about the typical Stern-Gerlach (SG) process for a spin $1/2$ atom, we ask what happens if a macroscopic body consisting of many (e.g. $N\sim\mathcal{O}(10^{23}))$ spins is set under an inhomogeneous magnetic field of a SG set-up. In general, with many spins oriented in different directions, the problem does not arise, as the bound-state Hamiltonian does not allow each component particle to split à la SG into spatially separated sub-wavepackets. Logically, however, one cannot exclude the special situation in which a large number of spins are all oriented in the same direction. Such a large spin system might continue to behave quantum-mechanically, e.g. with a typical SG spreading of the wavepackets, even in the $S/\hbar\to\infty$ limit.
It will be shown that the state $|S,{\mathbf{n}}\rangle$, the state in which the large spin is oriented in some direction ${\mathbf{n}}$, behaves instead as a classical angular momentum, yielding the unique trajectory in the SG processes (Sec. 2). This result leads us to consider more carefully in Sec. 3 the relation between the space of quantum-mechanical angular-momentum states and that of classical angular momenta, and to identify the special class of quantum angular momentum states to be put in correspondence with the classical angular momentum (see Eqs. (3.11), (3.12), (3.13) below). The content of the present work is to check the consistency of this picture, through the analysis of the SG experiments (Sec. 2), the angular-momentum composition rule (Sec. 4), and the rotation matrix (Sec. 5). We conclude with a brief reflection in Sec. 6.
2 Stern-Gerlach experiments
2.1 Spin $1/2$ atom
For definiteness let us take an incident atom with spin $S=\tfrac{1}{2}$ directed along a definite, but generic, direction ${\bf n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, i.e.
$\Psi=\psi({\bf r},t)\,|{\bf n}\rangle\;,$ (2.1)
When the atom passes through a region with an inhomogeneous magnetic field,
${\mathbf{B}}=(0,0,B(z)),\qquad\frac{\partial B(z)}{\partial z}\neq 0\;,$ (2.3)
the wave packet splits into two subpackets, with respective weights, $|c_{1}|^{2}$ and $|c_{2}|^{2}$: they manifest as the relative intensities of the image bands on the screen at the end.
The original Stern-Gerlach experiment [3] has had a fundamental impact to the development of quantum theory, in proving the quantization of the angular momentum, the existence of half-integer spin, and more generally, demonstrating the quantum-mechanical nature of the center of mass (CM) of the silver atom [2].
2.2 Large spin
A classical particle with a magnetic moment directed towards
@@ 2.2 Large spin
moves in an inhomogeneous magnetic field (2.3), according to Newton’s equations,
$m\,{\dot{\bf r}}={\bf p}\;,\qquad\frac{dp_{x}}{dt}=\frac{dp_{y}}{dt}=0\;,\quad\frac{dp_{z}}{dt}=F_{z}=\frac{\partial}{\partial z}{\boldsymbol{\mu}}\cdot{\bf B}\;.$ (2.5)
The way the unique trajectory for a classical particle emerges from quantum mechanics has been discussed in Ref. [1], where the magnetic moment is the expectation value
@@
taken in the internal bound-state wave function $|\Psi\rangle$ where $\mu_{i}$ and $\frac{e_{i}{\bf\ell}_{i}}{2m_{i}c}$ denote the intrinsic magnetic moment and the one due to the orbital motion of the $i$-th constituent atom or molecule ($i=1,2,\ldots,N$), respectively. Clearly, in general, the considerations made in Sec. 2.1 for a spin $1/2$ atom, with a doubly split wavepacket, cannot be simply generalized to (or compared with) a classical body (2.6) with $N\sim\mathcal{O}(10^{23})$ constituents.
@@
Typically, the spins are oriented along different directions, but the particles carrying them are bound in the macroscopic body made of atomic, molecular and larger structures. The wave function of particles in bound states does not split in sub-wavepackets, as that of an isolated spin $\tfrac{1}{2}$ atom does, under an inhomogeneous magnetic field. The bound-state Hamiltonian does not allow it  [^2] This discussion is indeed closely parallel to the question of quantum diffusion. Unlike free particles, particles in bound states (the electrons in atoms; atoms in molecules, etc.) do not diffuse, as they move in binding potentials. This is one of the elements for the emergence of the classical mechanics, with unique trajectories for macroscopic bodies. As for the center-of-mass (CM) wave function of an isolated macroscopic body, its free quantum diffusion is suppressed by mass.. What determines the motion of the CM of the body is the average magnetic moment, (2.6). Logically, nevertheless, one cannot exclude special cases (e.g. a magnetized piece of metal) of all spins of the body oriented in the same direction. One might wonder how the wave function of such a particle with spin $S\to\infty$ behaves, under an inhomogeneous magnetic field.
@@
The question is whether the conditions enlisted in Ref. [1] (see Appendix A), for the emergence of classical mechanics for a macroscopic body with a unique trajectory, are sufficient, or whether some extra condition (or a new mechanism) is required, to suppress possible wide spreading of the wave function into many sub-packets (see Fig. 1 for a small spin, $S=13/2$) under an inhomogeneous magnetic field.
@@
Figure 1: The distribution $|c_{k}|^{2}$ in $k$, i.e. in possible values of $S_{z}$, $-S\leq S_{z}\leq S$, for a spin $S=13/2$ particle in the state (2.7), with $\theta=\pi/2$ (center, orange dots) and $\theta=\pi/4$ (right, blue dots).
The answer turns out to be simple but somewhat unexpected. The following discussion, up to Eq. (2.15), is taken almost verbatim from Sec. 3.1.3 of Ref. [2]; for the present work it represents the starting point, and serves as the main tool of analysis below.
Consider the state of a spin $S$ directed towards a direction ${\mathbf{n}}$,
$\left({\hat{\mathbf{S}}}\cdot{\mathbf{n}}\right)\,|S,{\mathbf{n}}\rangle=S\,|S,{\mathbf{n}}\rangle\;,\qquad{\bf n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)\;.$ (2.7)
Choosing the magnetic field (and its gradients) in the ${z}$-direction, we need to express $|S,{\mathbf{n}}\rangle$ as a superposition of the eigenstates of ${\hat{S_{z}}}$ ($S=N/2$),
$|S,{\mathbf{n}}\rangle=\sum_{k=0}^{N}c_{k}\,|S,S_{z}=M\rangle\;,\qquad M=-\frac{N}{2}+k\;,\quad(k=0,1,\ldots,N)\;.$ (2.8)
The expansion coefficients $c_{k}$ are known:
$c_{k}={\binom{N}{k}}^{1/2}\,e^{-iM\phi/2}\left(\cos\tfrac{\theta}{2}\right)^{k}\left(\sin\tfrac{\theta}{2}\right)^{N-k}\;,\qquad\sum_{k=0}^{N}|c_{k}|^{2}=1\;,$ (2.9)
where $\binom{N}{k}=N!/k!(N-k)!$ are the binomial coefficients.
To get (2.9), let us consider (2.7) as a direct product of $N$ spin $\tfrac{1}{2}$ particles, (2.2),
$|S,{\mathbf{n}}\rangle=|{\bf n}\rangle_{1}\otimes|{\bf n}\rangle_{2}\otimes\cdots|{\bf n}\rangle_{N}\;,$ (2.10)
all oriented in the same direction. Expanding and collecting terms with a fixed $k$ (the number of spin-up particles) one arrives at Eq. (2.9).
Using Stirling’s formula, one can derive, for large values of both $N$ and $k$ with $x=k/N$ fixed, the distribution across different values of $S_{z}=M$,
$|c_{k}|^{2}\simeq e^{Nf(x)}\;,\qquad x=k/N\;,$ (2.11)
with
The saddle-point approximation, valid at $N\to\infty$, yields
and thus
@@
$S_{z}=M=N(x-\tfrac{1}{2})=S\,(2\cos^{2}\tfrac{\theta}{2}-1)=S\,\cos\theta\;.$ (2.15)
This means that a large spin ($S\gg\hbar$) quantum particle with spin directed towards ${\mathbf{n}}$, in a Stern-Gerlach setting with an inhomogeneous magnetic field (2.3), moves along a single trajectory of a classical particle with $S_{z}=S\,\cos\theta$, instead of spreading over a wide range of split sub-packet trajectories covering $-S\leq S_{z}\leq S$. See Fig. 2, cfr. Fig. 1.
Figure 2: The distribution $|c_{k}|^{2}$ in $k$, i.e. in possible values of $S_{z}$, for a spin $S=1\cdot 10^{3}$ particle in the state (2.7), with $\theta=\pi/2$ (center, orange peak) and $\theta=\pi/4$ (right, blue peak). The particle starts looking like classical.
This (perhaps) somewhat surprising result shows that quantum mechanics (QM) takes care of itself, so to speak, in ensuring that a large spin particle ($S/\hbar\to\infty$) traces a classical trajectory. This is consistent with the known general behavior of the wave function in the semi-classical limit ($\hbar\to 0$)  [^3] Of course this does not mean that the classical limit necessarily requires or implies $S/\hbar\to\infty$, though.. Note that if the value $S\gg\hbar$ is understood as due to the large number of spin $1/2$ particles composing it (see the comment after (2.9)), the spike (2.13), (2.14) can be understood as due to the accumulation of a huge number of microstates giving $S_{z}\simeq M$.
Another, equivalent way to see the shrinkage of the distribution in $S_{z}$, (2.14), is to study its dispersion. Given the relative weights $|c_{k}|^{2}$, (2.9), where $S_{z}=-S+k$ and $S=\tfrac{N}{2}$, with $k=0,1,\ldots,N$, it is a simple exercise to find that
$\langle S_{z}\rangle=S\,\cos\theta\;,\qquad\langle S_{z}^{2}\rangle=S^{2}\cos^{2}\theta+\frac{S}{2}\sin^{2}\theta\;,$ (2.16)
so that the dispersion (the fluctuation width) is given by
$(\Delta S_{z})^{2}=\langle(S_{z}-\langle S_{z}\rangle)^{2}\rangle=\langle S_{z}^{2}\rangle-\langle S_{z}\rangle^{2}=\frac{S}{2}\sin^{2}\theta\;.$ (2.17)
It follows that
$\lim_{S/\hbar\to\infty}\frac{|\Delta S_{z}|}{S}=0\;,$ (2.18)
i.e. it is an infinitely narrow Gaussian. Note that the SG set-up is indeed a measurement of $S_{z}$ in the state (2.7). The fact that the result is always $S_{z}\sim\langle S_{z}\rangle=S\,\cos\theta$ with no dispersion, means that it is a classical angular momentum $\sim S{\mathbf{n}}$ (see Eqs. (3.11), (3.12) below).
2.2.1 Density matrix
The same phenomenon can be seen also by using the density matrix. Being the state (2.7) a pure state, its $(2S+1)\times(2S+1)$ density matrix is given by
$\rho_{k\ell}=c_{k}c_{\ell}^{*}\,,$ (2.19)
where $c_{k}(c_{\ell})$ are given in Eq. (2.9). In the limit as $S\to\infty$, it can be observed that it essentially reduces to a single diagonal element, provided that the fluctuations of the order of $\sqrt{S}$ are neglected in comparison to $S$. The density matrix ${\boldsymbol{\rho}}$ for $S=500$, with $(\theta,\phi)=(\frac{\pi}{2},0)$ and $(\theta,\phi)=(\frac{\pi}{4},0)$, is depicted in Figs. 3.
Figure 3: The density matrix ${\boldsymbol{\rho}}$ for the state (2.7), of spin $S=5\cdot 10^{2}$, with ($a$): $(\theta,\phi)=(\tfrac{\pi}{2},0)$ and ($b$): $(\theta,\phi)=(\tfrac{\pi}{4},0)$. The height of the plots indicates the matrix elements $\rho_{k\ell}$. Note that ${\boldsymbol{\rho}}$ is real for $\phi=0$.
2.2.2 The state preparation and successive SG set-ups
For a small $S$, as it is well known, the SG set-up can be used to prepare any desired state $|S,S_{z}=M\rangle$, by an intentionally biased SG apparatus, e.g. in which all the split sub-wavepackets are blocked, except for the one corresponding to $|S,S_{z}=M\rangle$. Equivalently, one may use the so-called null-measurements (see Ref. [4] for a recent review). In any case, for a small finite $S$, where there are $12S+$ 1 well-separated sub-packets (see Fig. 1), there is no difficulty in extracting any desired quantum state for subsequent experiments (state preparation).
In the large-spin limit ($S\to\infty$) we are considering, since there is essentially only one wavepacket present (see (2.15)), there is no possibility of extracting or selecting the desired sub-wavepacket. The experimental resolution would not allow it. We have in mind a macroscopic scenario where, in comparison to the finite region (in principle) occupied by the waves $-S\leq M\leq S$, the fluctuations of the values of $M$ around (2.15) are of the order of $\sqrt{S}\ll S$, i.e. of zero width.
We might consider introducing a second set of Stern-Gerlach (SG) magnetic fields directed in a different direction, denoted by ${\hat{z}}^{\prime}$, as a new quantization axis. This setup allows us to study the state emerging from the first set of SG magnetic fields (without the presence of an imaging screen),
@@ 2.2.2 The state preparation and successive SG set-ups
$|S,{\mathbf{n}}\rangle\sim c_{k}\,|S,S_{z}=M\rangle\;,\qquad k\sim 2S\cos^{2}\tfrac{\theta}{2}\;,\qquad M\sim S\,\cos\theta\;.$ (2.20)
Since the sum of states around $k\sim 2S\cos^{2}\left(\frac{\theta}{2}\right)$ is essentially equal to the complete sum, this state can be rewritten in terms of the eigenvalues $M^{\prime}=-S+k^{\prime}$, where $k^{\prime}=0,1,\ldots 2S$ correspond to the eigenvalues of $S_{z^{\prime}}$ as:
$|S,{\mathbf{n}}\rangle$ $=$ $\sum_{k,k^{\prime}}c_{k}\,|S,M^{\prime}\rangle\langle S,M^{\prime}|S,M\rangle$ (2.21)
$=$ $\sum_{k,k^{\prime}}|S,M^{\prime}\rangle\langle S,M^{\prime}|S,M\rangle\langle S,M|S,{\mathbf{n}}\rangle$
$=$ $\sum_{k^{\prime}}|S,M^{\prime}\rangle\langle S,M^{\prime}|S,{\mathbf{n}}\rangle=\sum_{k^{\prime}}c_{k^{\prime}}|S,M^{\prime}\rangle\;.$
This represents an expansion of the original state $|S,{\mathbf{n}}\rangle$ as a linear combination of the eigenstates of $S_{z^{\prime}}$. In the limit $S\to\infty$, the results from Section 2.2 tell us that the sum is dominated by the small region around $S_{z^{\prime}}\sim S,\cos({\mathbf{n}}\cdot{\hat{z}}^{\prime})$, with negligible fluctuations. The trajectory remains unique, regardless of the choice of quantization axis (i.e. the direction of the magnetic field) ${\hat{z}}^{\prime}$.
2.2.3 States almost directed towards ${\vec{\mathbf{n}}}$
Obviously, in the large spin limit, other states close to $|S,{\mathbf{n}}\rangle$,
$\left({\hat{\mathbf{S}}}\cdot{\mathbf{n}}\right)\,|S,{\mathbf{n}},q\rangle=(S-q)\,|S,{\mathbf{n}},q\rangle\;,\qquad q=1,2,\ldots\,\,\,{\rm fixed}$ (2.22)
$|S,{\mathbf{n}},1\rangle$ $=$ $\sum_{\ell=0}^{N-1}\frac{1}{\sqrt{N}}\,|{\bf n}\rangle_{1}\otimes\cdots|{\bf n}\rangle_{\ell}\otimes|-{\bf n}\rangle\otimes|{\bf n}\rangle_{\ell+2}\otimes\cdots|{\bf n}\rangle_{N}\;,$ (2.23)
$=$ $\sum_{\ell=0}^{N-1}\frac{1}{\sqrt{N}}|S_{1},{\mathbf{n}}\rangle\otimes|-{\bf n}\rangle\otimes|S_{2},{\mathbf{n}}\rangle\;,$
where $|S_{1},{\mathbf{n}}\rangle$ and $|S_{2},{\mathbf{n}}\rangle$ ($S_{1}=\tfrac{\ell}{2}\;,S_{2}=\tfrac{N-\ell-1}{2}$), are two spins oriented towards ${\mathbf{n}}$, (2.8), as already studied in Sec. 2.2. In the $S\to\infty$ limit, the sum is dominated by terms with $\ell\sim\mathcal{O}(N)\to\infty$. The SG projection of this state onto the $S_{z}$ eigenstates is therefore approximated by
$S_{z}\sim S_{1\,z}+S_{2\,z}=S_{1}\cos\theta+S_{2}\cos\theta\simeq S\,\cos\theta\;,$ (2.24)
where the results of Sec. 2.2 have been used. A similar result holds for $|S,{\mathbf{n}},q\rangle$ for any finite $q$, showing that all these states behave as a classical angular momentum, $S\,{\mathbf{n}}$.
3 Quantum and classical angular momentum states
Before proceeding, we must discuss the correspondence between the quantum mechanical angular momentum states and their classical counterparts, more carefully. The spin $1/2$ (two-state system) wave function is given by
@@ 3 Quantum and classical angular momentum states
where $(c_{1},c_{2})\sim\lambda(c_{1},c_{2})$ with $\lambda\in{\mathbf{C}}/\{0\}$, describes the points of $\mathbf{CP}^{1}\sim S^{2}$. We note that for a spin $1/2$, any pure state (3.1) can be interpreted as the state in which the spin is oriented in some space direction ${\mathbf{n}}$, as in Eq. (2.2), without loss of generality. This is because the manifold of pure spin $1/2$ quantum states and that of the unit space vector are both $S^{2}\sim\mathbf{CP}^{1}$. Indeed, given the state (2.2), there is always a rotation matrix such that
$D^{1/2}(0,\theta,\phi)\cdot\left(\begin{array}[]{c}e^{-i\phi/2}\cos\tfrac{\theta}{2}\\ e^{i\phi/2}\sin\tfrac{\theta}{2}\end{array}\right)=\left(\begin{array}[]{c}1\\ 0\end{array}\right)\;.$ (3.2)
@@
We recall that the general rotation matrix for a spin $1/2$ is given, using Euler angles  [^4] Note that the third Euler angle (the rotation angle $\gamma$ about the final $z$-axis) is redundant here, as it gives only the phase $e^{i\gamma/2}$, and has been set to 00 in (3.2). ($\gamma,\theta,\phi$), as follows
@@
The situation is different for a generic spin $j=N/2$  [^5] In this section we denote the general angular momentum (including orbital angular momentum) by $j$, rather than $S$, a symbol adopted commonly to indicate the spin (the intrinsic angular momentum of a particle). In this work we will use both $j$ and $S$. Likewise, we shall use the terms “spin” and “angular momentum”, interchangeably.. The wave function now has the form,
$|\psi\rangle=\sum_{k=0}^{N}b_{k}|j,j_{z}=-j+k\rangle$ (3.4)
@@
$(b_{0},b_{1},\ldots,b_{N})\sim\lambda(b_{0},b_{1},\ldots,b_{N})\;,\qquad\lambda\in{\mathbf{C}}/{(0)}\;,$ (3.5)
are the coordinates of the space $\mathbf{CP}^{N}$. On the other hand, the variety of the directions in the three-dimensional space, ${\mathbf{n}}$, is always $S^{2}\sim\mathbf{CP}^{1}$. It means that the state of the spin oriented towards a definite direction (2.8), is not a general pure spin $j$ state, but belongs to a small subset, $\mathbf{CP}^{1}\subset\mathbf{CP}^{2j}$. In other words, not all states of spin $j$, as described in (3.4), can be transformed by a rotation matrix (selecting an appropriate new ${\hat{z}}$ axis) into the form:
$(1,0,\ldots,0)^{T}\;.$ (3.6)
These considerations highlight a potential subtlety in our argument that quantum mechanical angular momentum states somehow become classical in the large $j$ limit.
Let us recall that because of the commutation relations
@@
In the state, $|j,m\rangle$, where ${\hat{J}}_{z}$ has a definite value, $m\,\hbar$, ${\hat{J}}_{x}$ and ${\hat{J}}_{y}$ are fluctuating. Their magnitudes $\Delta{J}_{x,y}=\langle j,m|{\hat{J}}_{x,y}^{2}|j,m\rangle^{1/2}$ are found to be
@@
Therefore, in a generic state $|j,m\rangle$, there are strong quantum fluctuations of $J_{x}$ and $J_{y}$ at $j\to\infty$. There is no way such a state can be associated with a classical angular momentum, which requires all three components to be well-defined simultaneously.
The exception occurs for states $|j,\pm j\rangle$, for which

### v3

@@ Abstract
Quantum states of a spin $\tfrac{1}{2}$ (a qubit) are parametrized by the space ${\mathbf{CP}}^{1}\sim S^{2}$, the Bloch sphere. A spin $j$ for a generic $j$ (a $12j+$ 1-state system) is represented instead by a point of a larger space, ${\mathbf{CP}}^{2j}$. Here we study the state of a single angular momentum/spin in the limit, $j\to\infty$. The special class of states $|j,{\mathbf{n}}\rangle\in{\mathbf{CP}}^{2j}$, with spin oriented towards definite spatial directions ${\mathbf{n}}\in S^{2}$, i.e., $({\hat{\mathbf{J}}}\cdot{\mathbf{n}})\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle$, are found to behave as classical angular momenta, $j\,{\mathbf{n}}$, in this limit. Vice versa, general spin states in ${\mathbf{CP}}^{2j}$ do not become classical, even at large $j$. We discuss these questions, by analysing the Stern-Gerlach processes, the angular-momentum composition rule, and the rotation matrix. Our observations help to clarify better how classical mechanics emerges from quantum mechanics in this context (e.g., with unique trajectories for a particle carrying a large spin), and to make the widespread idea that large spins somehow become classical, more precise.
@@ 1 Introduction
It is a widely shared view that in the limit of large spin (angular momentum) a quantum-mechanical spin somehow becomes classical [1]-[21]. After all, the large spin limit ($j/\hbar\to\infty$) is equivalent to the $\hbar\to 0$ ($j$  fixed) limit, and, in accordance with the general idea of the semi-classical limit in quantum mechanics, a large-angular-momentum state might be expected to behave as a classical angular momentum. But to the best of the authors’ knowledge, it has never been clearly shown whether this indeed occurs or if it does, exactly how. It is the aim of this note to help filling in this gap.
The question was put under a sharp focus in a general discussion on emergence of classical mechanics for a macroscopic body from quantum mechanics, as formulated in [22, 23]. Knowing about the typical Stern-Gerlach (SG) process for a spin $\frac{1}{2}$ atom, one asks what happens if a macroscopic body consisting of many (e.g. $N\sim\mathcal{O}(10^{23}))$ spins is set under an inhomogeneous magnetic field of a SG set-up. A classical particle with a magnetic moment directed towards
@@
moves in an inhomogeneous magnetic field according to Newton’s equations,
$m\,{\dot{\bf r}}={\bf p}\;,\qquad\frac{d{\bf p}}{dt}={\bf F}=\nabla({\boldsymbol{\mu}}\cdot{\bf B})\;.$ (1.2)
The way how such a unique trajectory for a classical particle emerges from quantum mechanics has been discussed in Ref. [22]. The magnetic moment of a macroscopic particle is the expectation value
@@
taken in the internal bound-state wave function $|\Psi\rangle$ where $\mu_{i}$ and $\frac{e_{i}{\bf\ell}_{i}}{2m_{i}c}$ denote the intrinsic magnetic moment and the one due to the orbital motion of the $i$-th constituent atom or molecule ($i=1,2,\ldots,N$), respectively. Clearly, in general, the well-known picture for a spin $\frac{1}{2}$ atom, with a doubly split wavepacket, does not generalize to a classical body (1.3) with $N\sim\mathcal{O}(10^{23})$ constituents.
@@
Typically, the spins are oriented along different directions, but the crucial point is that the component particles carrying them are bound in the macroscopic body in various atomic, molecular or crystalline structures. The wave function of particles in bound states does not split in sub-wavepackets, in contrast to that of an isolated spin $\tfrac{1}{2}$ atom, under an inhomogeneous magnetic field. The bound-state Hamiltonian does not allow it  [^1] This discussion is closely parallel to the question of quantum diffusion. Unlike free particles, particles in bound states (the electrons in atoms; atoms in molecules, etc.) do not diffuse, as they move in binding potentials. This is one of the elements for the emergence of the classical mechanics, with unique trajectories for macroscopic bodies. As for the center-of-mass (CM) wave function of an isolated macroscopic body, its free quantum diffusion is simply suppressed by mass.. What determines the motion of the CM of the body is the average magnetic moment, (1.3).
@@
Logically, however, one cannot exclude special cases (e.g. a magnetized piece of metal) with all spins inside the body oriented in the same direction. One might wonder how the wave function of such a particle with spin $j\to\infty$ behaves, under an inhomogeneous magnetic field. The question is whether the three conditions recognized in [22] for the emergence of classical mechanics for a macroscopic body with a unique trajectory, reviewed in Appendix A here, are indeed sufficient. Or, is some extra condition, or a new unknown mechanism, needed, to suppress possible wide spreading of the wave function into many sub-packets (such as in Fig. 1 for a small spin, $j=13/2$) under an inhomogeneous magnetic field?
@@
Figure 1: The distribution $|c_{k}|^{2}$ in $k$, i.e. in possible values of $J_{z}=m$, $-j\leq m\leq j$ ($m=-j+k$) for a spin $j=13/2$ particle in the state (2.26), (2.27), with $\theta=\pi/2$ (center, orange dots) or with $\theta=\pi/4$ (right, blue dots).
The aim of this work is to answer these questions. We shall first discuss in Sec. 2 some mathematical and physical aspects of the quantum states of a small and large angular momenta, and compare them with the properties of a classical angular momentum. Considerations of quantum fluctuations of various angular momentum states lead us to propose that a particular class of states, $|j,{\mathbf{n}}\rangle$,
${\hat{\mathbf{J}}}^{2}\,|j,{\mathbf{n}}\rangle=j(j+1)|j,{\mathbf{n}}\rangle\,,\qquad({\hat{\mathbf{J}}}\cdot{\mathbf{n}})\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle\,,$ (1.4)
where ${\mathbf{n}}$ is a unit vector directed towards $(\theta,\phi)$ direction,
${\mathbf{n}}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)\;.$ (1.5)
are to be identified with a classical angular momentum, $j\,{\mathbf{n}}$, in the large $j$ limit. The states (1.4) are known in the literature as coherent spin states or Bloch states [5, 10, 6, 11, 7].
Vice versa, generic states of large spin will be found to remain quantum mechanical, with large fluctuations, even in the limit, $j\to\infty$, showing that the often stated belief that a large spin (angular momentum) becomes classical, should not be taken for granted, literally.
In subsequent sections, we are going to examine these ideas through the analysis of the Stern-Gerlach processes (Sec. 3), the orbital angular momenta (Sec. 4), the angular-momentum addition rule (Sec. 5), and the rotation matrix (Sec. 6). Generic spin $j$ states far from (1.4) are shown to remain quantum mechanical even in the limit $j\to\infty$: the fate of these states will be discussed in Sec. 7. Conclusion and a few more general reflections are in Sec. 8.
2 Space of quantum spin states and classical angular momenta
The generic pure spin $\frac{1}{2}$ (two-state system) state is described by the wave function
where
$(c_{1},c_{2})\sim\lambda(c_{1},c_{2})\;,\qquad\lambda\in{\mathbf{C}}/\{0\}\,$ (2.2)
@@ 2 Space of quantum spin states and classical angular momenta
${\mathbf{n}}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)\;,$ (2.3)
that is
${\hat{J}}^{2}\left|\frac{1}{2},{\mathbf{n}}\right\rangle=\frac{3}{4}\left|\frac{1}{2},{\mathbf{n}}\right\rangle\;,\qquad\left({\hat{\mathbf{J}}}\cdot{\mathbf{n}}\right)\,\left|\frac{1}{2},{\mathbf{n}}\right\rangle=\frac{1}{2}\,\left|\frac{1}{2},{\mathbf{n}}\right\rangle\;,$ (2.4)
@@
$|{\bf n}\rangle\equiv\left|\frac{1}{2},{\mathbf{n}}\right\rangle=c_{1}|\!\uparrow\rangle+c_{2}|\!\downarrow\rangle\;,\qquad c_{1}=e^{-i\phi/2}\cos\tfrac{\theta}{2}\;,\qquad c_{2}=e^{i\phi/2}\sin\tfrac{\theta}{2}\;.$ (2.5)
@@
without loss of generality. This is so because the space of pure spin $\frac{1}{2}$ quantum states and that of the unit space vector coincide: they are both $S^{2}\sim\mathbf{CP}^{1}$. Indeed, given the state (2.5), there is always a rotation matrix such that
$D^{1/2}(0,\theta,\phi)\cdot\left(\begin{array}[]{c}e^{-i\phi/2}\cos\tfrac{\theta}{2}\\ e^{i\phi/2}\sin\tfrac{\theta}{2}\end{array}\right)=\left(\begin{array}[]{c}1\\ 0\end{array}\right)\;,$ (2.6)
@@
where the general rotation matrix for a spin $\frac{1}{2}$ is given, using Euler angles  [^2] Note that the third Euler angle (the rotation angle $\gamma$ about the final $z$-axis) is redundant here, as it gives only the phase $e^{i\gamma/2}$, and has been set to 00 in Eq.(2.6). ($\gamma,\theta,\phi$), by
@@
The fact that any spin $\frac{1}{2}$ quantum state can be associated with a definite space direction, however, does not mean that it can be identified with a classical angular momentum, $\frac{1}{2}\hbar{\mathbf{n}}$, as the fluctuations in the directions transverse to ${\mathbf{n}}$ are always of the same order of the spin magnitude itself (see Eq.(2.13) below, with $j=\frac{1}{2}$, $m=\pm\frac{1}{2}$). It is always a quantum mechanical system.
Let us now consider a generic spin $j$ state. It is a special type of $12j+$ 1-state quantum system. Its general wave function has the form,
$|\psi\rangle=\sum_{m=-j}^{j}b_{m}|j,j_{z}=m\rangle$ (2.8)
@@
$(b_{-j},b_{-j+1},\ldots,b_{j})\sim\lambda(b_{-j},b_{-j+1},\ldots,b_{j})\;,\qquad\lambda\in{\mathbf{C}}/{(0)}\;,$ (2.9)
are the coordinates of the complex projective space $\mathbf{CP}^{2j}$ 333 As is well known, the manifold of pure quantum states of any $n$-state system is $\mathbf{CP}^{n-1}$ [24]. The spin states (2.9) are special cases with $n=2j+1$. . On the other hand, the variety of the directions in the three-dimensional space, ${\mathbf{n}}$, is always $S^{2}\sim\mathbf{CP}^{1}$. It means that this time not all states of spin $j$ state, (2.8), can be transformed by a rotation matrix (selecting an appropriate new ${\hat{z}}$ axis) into the form:
$(b_{-j}^{\prime},b_{-j+1}^{\prime},\ldots,b_{j}^{\prime})\sim(0,0,\ldots,0,1)\;.$ (2.10)
This observation is the first hint that there are some subtleties in the way classical picture of angular momentum emerges from quantum mechanics at large $j$. It is the purpose of the present note to elaborate on this point.
Let us start with the basic properties of quantum mechanical angular momentum. These are of course well-known textbook materials. Because of the commutation relations
@@
In the state, $|j,m\rangle$, where ${\hat{J}}_{z}$ has a definite value, $m\,\hbar$, ${\hat{J}}_{x}$ and ${\hat{J}}_{y}$ are fluctuating. Their magnitudes $\Delta{J}_{x,y}=\langle j,m|{\hat{J}}_{x,y}^{2}|j,m\rangle^{1/2}$ are given by
@@
Namely, in a generic state $|j,m\rangle$, there are strong quantum fluctuations of $J_{x}$ and $J_{y}$ at $j\to\infty$, of the same order of magnitude with $j$ itself. There is no way such a state can be associated with a classical angular momentum, which requires all three components to be well-defined simultaneously. The exception occurs for states $|j,\pm j\rangle$, for which